# Correlators are simpler than wavefunctions

### v2

@@ I Introduction
$\Psi[\Phi]$ $=\int_{\varphi(t=-\infty)}^{\varphi(t=0)}\mathcal{D}\varphi\>e^{-i\int_{-\infty}^{0}dt\,\int_{-\infty}^{+\infty}d^{n}x\,S[\varphi]},$ (1)
@@
$\langle\Phi^{N}\rangle$ $=\int_{\varphi(t=-\infty)}^{\varphi(t=\infty)}\mathcal{D}\varphi\>\prod_{i}^{N}\Phi(0,\vec{x}_{i})\>|\Psi[\varphi]|^{2}.$ (2)
@@
Both the wavefunction and the correlator admit expansion in terms of Feynman graphs, examples of which are shown in figure 1. We will focus most of our analysis at this graph level. For a graph $G$, with vertex parameters $x_{v}$ and edge parameters $y_{e}$, the wavefunction contribution takes the form:
@@
$\langle G\rangle=\int_{-\infty}^{\infty}dt^{V}e^{-\sum_{v}x_{v}|t_{v}|}\prod_{e}G_{F}(t_{e},t^{\prime}_{e};y_{e}).$ (4)
@@
At the level of individual graphs, there is growing evidence to support this philosophical hope for simplification. First, correlators exhibit only a subset of the singularities present in the wavefunction [38], drastically reducing the complexity of the relevant function space [39, 40]. Second, the ‘melonic’ loop integrands behave much like their tree-level counterparts [37]. A key structural reason for these simplifications is that the wavefunction is computed over a half-time domain, from the asymptotic past up to the boundary. In contrast, the correlator arises from integral over the full time domain  [^1] The equivalence between in-in formalism and in-out correlation functions has been discussed in [45] for flat space, and [37] for de Sitter correlators..
@@
Systematic expansion—Although the factorisation properties of the wavefunction and correlator are closely related, we show that expansions around a given pole are systematically simpler for correlators. In particular, we verify that for the correlator, the first sub-leading term always vanishes. Furthermore, for the total-energy pole, the next two subleading terms take remarkably compact forms that can be generated by acting with a simple differential operator on the amplitude.
@@ II Wavefunction vs correlator
Wavefunction singularities—The wavefunction develops singularities whenever the total energy entering any connected sub-graph $\mathfrak{g}\subseteq G$ vanishes [1]. In this context, we use the convention that any edge entering the sub-graph from outside, is replaced by a half-edge attached to the corresponding vertex, with the half-edge inheriting the edge variable of the original. For instance, consider the sub-graph of the two-cycle shown in figure 1 consisting of only the top edge, for which we have:
@@
As an example, the singularities for the bubble integrand are given by
@@
This absence of poles is seen directly from the integral representation (4). It should be clear that any $2y_{e}$ appearing in a pole can only arise because a step function $\theta(t_{j}-t_{i})$ forces the integration of a given $t_{i}$ to have a boundary at $t_{j}$. To prove that this cannot happen, we need only expand each propagator of (4) in terms of step functions. Then, we can pick the terms of the integral that isolate the step-function $\theta(t_{j}-t_{i})$. There will be two terms (out of a sum of four) that could be problematic, one is:
$\int_{-\infty}^{\infty}dt_{i}dt_{j}e^{x_{i}t_{i}}e^{-x_{j}t_{j}}e^{y(t_{i}-t_{j})}\theta(t_{i}-t_{j})\theta(-t_{i})\theta(t_{j})\,$, (14)
which is zero, due to the incompatibility of the step functions. The other part is:
$\int_{-\infty}^{\infty}dt_{i}dt_{j}e^{x_{i}t_{i}}e^{-x_{j}t_{j}}e^{y(t_{i}-t_{j})}\theta(t_{i}-t_{j})\theta(t_{i})\theta(-t_{j})\,$, (15)
@@
This allows us to restrict our analysis of the correalator to graphs in which each pair of vertices is connected by at most one edge.
@@ III Soft Limits
Connected subgraphs—We now explore the case where we take all vertex variables in a graph $G$ to zero, except those in a connected subgraph $\mathfrak{g}$. Choosing $x_{v_{*}}$, where $v_{*}\in\mathfrak{g}$, as the root in (19), then the integral factorizes into an integral over the variables in $\mathfrak{g}$, and an integral over the edges in $G/\mathfrak{g}$:
$\langle\widehat{G}\rangle$ $=\int\limits_{-\infty}^{\infty}dt_{*}ds_{e^{\prime}}\prod_{e^{\prime}\in E(\mathfrak{g})}\widehat{G}_{F}(s_{e^{\prime}})\int\limits_{-\infty}^{\infty}ds_{e}\prod_{e\in E(G/\mathfrak{g})}\widehat{G}_{F}(s_{e}).$
$=\langle\hat{\mathfrak{g}}\rangle\times\mathcal{A}({G/\mathfrak{g}})\,.$ (20)
In this limit the correlator factorizes into the correlator of $\mathfrak{g}$, and the amplitude of a graph where $\mathfrak{g}$ is contracted to a single vertex, which we denote by $G/\mathfrak{g}$. A particularly interesting case is when $\mathfrak{g}$ itself is a single vertex, then the property holds for loop graphs as well, and we obtain the $\ell_{0}$-integrated amplitude (and setting all the zeroth component of four-momenta to zero):
$\langle\widehat{G}\rangle|_{x_{v_{*}}\neq 0}=\frac{2}{x_{v_{*}}}\times\mathcal{A}(G),$ (21)
For a connected graph, the $l_{0}$-integrated amplitude $\mathcal{A}(G)$ satisfies a Berends-Giele type recursion
$\mathcal{A}(G)=\sum_{v\in V(G)}\prod_{\mathfrak{g}\in\kappa(G\setminus v)}\left.\frac{\mathcal{A}(\mathfrak{g})}{B(\mathfrak{g})}\right|_{x_{v}=0},$ (22)
where the product runs over the connected components of the graph $G\setminus v$. The recursion terminates on graphs consisting of a single vertex, for which we have $A(v)=1$.
@@
$\langle\widehat{G}\rangle=\left(\int\limits_{-\infty}^{\infty}ds_{e^{\prime}}\prod_{e\in E(G/P_{*})}\widehat{G}_{F}(s_{e^{\prime}})\right)\times$
@@
$\langle\widehat{G}\rangle=\left(\prod_{e\in E(G/P_{*})}\frac{1}{y_{e}}\right)\times$
$\sum_{p_{*}\in\{x_{*},x,y_{e}\}}\mathop{\mathrm{Res}}_{p\to ip_{*}}\frac{2x_{*}x\prod_{e\in E(P_{*})}2y_{e}}{(p^{2}+x_{*}^{2})(p^{2}+x^{2})\prod_{e\in E(P_{*})}(p^{2}+y_{e}^{2})}.$ (24)
@@ IV Expanding the correlator
We now analyze the simplifications that occur in expanding the correlator about an arbitrary pole, and present a systematic expansion in the total energy.
Tree-level $\mathcal{O}(\mathcal{E}_{G})$-expansion—We begin our analysis by examining the total-energy expansion of an arbitrary tree-level graph. To bring the integral into a form suitable for expansion we introduce the rescaled variables $x_{v}=\epsilon\bar{x}_{v}$ and $\tau=\epsilon t_{*}$ such that (19) becomes
$\langle G\rangle=\frac{1}{\epsilon}\int_{-\infty}^{\infty}d\tau\int_{-\infty}^{\infty}ds^{E}F(\epsilon)\prod_{e\in E(G)}G_{F}(s_{e}),$ (25)
@@
$\langle G\rangle^{(-1)}=\frac{1}{\epsilon}\int_{-\infty}^{\infty}d\tau F(0)\int_{-\infty}^{\infty}ds^{E}\prod_{e\in E(G)}G_{F}(s_{e}).$ (28)
@@
$\langle G\rangle^{(-1)}=\frac{2}{\sum_{v}x_{v}}\prod_{e\in G}\frac{1}{y_{e}^{2}}.$ (29)
@@
$\langle\widehat{G}\rangle^{(1)}=\frac{-\epsilon}{\sum_{v}x_{v}}\sum_{u\neq v}x_{u}x_{v}\int_{-\infty}^{\infty}ds^{E}\sum_{e\in P_{u}^{v}}s_{e}^{2}\prod_{vv^{\prime}\in E_{G}}\widehat{G}_{F}(s_{e}),$ (31)
@@
$\langle\widehat{G}\rangle^{(1)}=-\sum_{u\neq v}\frac{x_{u}x_{v}}{\mathcal{E}_{G}}\sum_{e\in P_{u}^{v}}\partial_{y_{e}}^{2}\prod_{e\in E(G)}\frac{2}{y_{e}},$ (32)
where the total energy is given by $\mathcal{E}_{G}=\sum_{v}x_{v}$.
@@
For an $n$-loop diagram $n$ variables $s_{e}$ will be constrained in terms of the others. Naturally, the equations (27) still hold true, and we recover the $\ell_{0}$-integrated amplitude in this limit for the $\mathcal{E}_{G}^{-1}$ coefficient, and the $\mathcal{E}_{G}^{0}$ coefficient also vanishes. However, for the $\mathcal{E}_{G}$ there is a slight complication. We find from (47), that the cross terms of the form $s_{e}s_{e^{\prime}}$ no longer vanish, given that the integral is no longer odd. This is simply because the argument of on of the Feynman propagators will be (33) as explained above. Therefore there is no obvious form for the differential operator acting on the amplitude which gives us the $\mathcal{E}_{G}$ term. Nevertheless, it is still possible to generate this coefficient. We first notice that, because of (33), we have a choice in how we parametrize the sum of $s_{e}$ in (47). This is simply the statement that for loop diagrams there is more than one path between two sites. Then, we can consider a one higher-loop graph where we added an edge between two sites, for this diagram we can parametrize (47) such that the path between these two sites is a single $s_{e}$, and thus we do not have the issue of cross terms not vanishing. Then, for loop graphs the differential operator is simply:
$\langle\widehat{G}\rangle^{(1)}=-\frac{1}{\sum_{v}x_{v}}\sum_{u\neq v}x_{u}x_{v}\partial_{y_{uv}}^{2}\mathcal{A}^{uv}(G)|_{y_{uv}\to 0},$ (34)
@@
$\langle\widehat{G}\rangle^{(2)}=-\frac{1}{\sum_{v}x_{v}}\sum_{u\neq v}x_{u}x_{v}\partial_{y_{uv}}^{3}\mathcal{A}^{uv}(G)|_{y_{uv}\to 0}.$ (36)
Beyond $\mathcal{O}(\mathcal{E}_{G}^{2})$ the coefficients become complicated polynomials of $x_{v}$, and it is no longer clear whether we can obtain them by acting with a differential operator on an amplitude.
@@
$x_{v}\rightarrow\epsilon x_{v}\text{ for all }v\in V(\mathfrak{g}),$
$y_{e}\rightarrow\epsilon y_{e}\text{ for all }e\in H(\mathfrak{g}),$ (37)
@@
$\langle G\rangle^{(-1)}=\frac{\mathcal{A}(\mathfrak{g})}{\mathcal{E}_{\mathfrak{g}}}\prod_{\mathfrak{h}\in\kappa(G\setminus\mathfrak{g})}\langle\mathfrak{h}\rangle,$ (38)
where the product is over all connected components of the graph $G$ with $\mathfrak{g}$ removed. Note, due to the limit we are taking, $\langle g\rangle$ is the correlator of the respective connected component with the edges connected to the central tube set to zero.
Furthermore, the first derivative with respect to $\epsilon$ is the same as in (27), therefore the sub-leading term in the series expansion around any pole of the correlator vanishes. This places strong constraints on the structure of the correlator. We have found in tree-level examples up to $4$ vertices—including both the $4$-chain and the star—that if we make an ansatz for the correlator inputing only the poles, demanding only the vanishing of subleading terms almost completely fixes the correlator.
@@ V $\text{Tr}(\phi^{3})$ correlator
Some of the results above naturally extend to the full correlator.
@@
$\langle\phi^{n}\rangle|_{|{\bf k}_{*}|\neq 0}=\frac{1}{|{\bf k}_{*}|}A_{n},$ (39)
@@
Pole expansion—The interesting structure in the expansion around any pole that we showed above for single graphs, generalizes for the sum of graphs. To obtain the expansion around a given pole we start by re-scaling every momenta (either external or internal), $\vec{k}_{i}\to\epsilon\,\vec{k}_{i}$, and then expand around $\epsilon=0$. The leading term will then be the usual factorization structure of the correlator, and the sub-leading term will vanish. For the total energy pole also the $\mathcal{O}(\mathcal{E}_{\text{tot}}^{1})$ term will have an interesting structure. We can reformulate our rule to obtain the operator acting on the amplitude to retrieve the $\mathcal{E}_{\text{tot}}$ term. The rule for (29) tells us that for every monomial $x_{i}\,x_{j}$ we should act with the sum over all double derivatives with respect to the every edge on the path from $x_{i}$ to $x_{j}$. To reformulate this as a full correlator rule it is best to think about it in terms of the momentum polygon for an $n$-point process. We know each triangulation of the polygon is dual to a graph, and the internal chords (parametrized by $|k_{ij}|$ of a triangulation are dual to edges (above parametrized by $y_{i}$) of the graph. Therefore, the rule of the $\mathcal{E}_{\text{tot}}$ operator is now that for each monomial $|\vec{k}_{i}|\,|\vec{k}_{j}|$ we take the sum over all the double derivatives with respect to the internal cords crossed when we draw a straight line from the outer edge $|\vec{k}_{i}|$ to the outer edge $|\vec{k}_{j}|$. This allows us to re-write (29) as:

### v3

@@ I Introduction
$\Psi[\Phi]$ $=\int_{\varphi(t=-\infty)}^{\varphi(t=0)=\Phi}\mathcal{D}\varphi\>e^{-i\int_{-\infty}^{0}dt\,\int_{-\infty}^{+\infty}d^{n}x\,S[\varphi]},$ (1)
@@
$\langle\Phi^{N}\rangle$ $=\int\mathcal{D}\Phi\>\prod_{i}^{N}\Phi(0,\vec{x}_{i})\>|\Psi[\Phi]|^{2}.$ (2)
@@
Both the wavefunction and the correlator admit expansion in terms of Feynman graphs, examples of which are shown in figure 1. We will focus most of our analysis at this graph level. For a graph $G$, with vertex parameters $x_{v}$ and edge parameters $y_{e}$, the wavefunction contribution takes the form (where we omit the loop integration):
@@
$\langle G\rangle=\int_{-\infty}^{\infty}dt^{V}\int e^{-\sum_{v}x_{v}|t_{v}|}\prod_{e}G_{F}(t_{e},t^{\prime}_{e};y_{e}).$ (4)
@@
At the level of individual graphs, there is growing evidence to support this philosophical hope for simplification. First, correlators exhibit only a subset of the singularities present in the wavefunction [38], drastically reducing the complexity of the relevant function space [39, 40]. Second, the ‘melonic’ loop integrands behave much like their tree-level counterparts [37]. A key structural reason for these simplifications is that the wavefunction is computed over a half-time domain, from the asymptotic past up to the boundary. In contrast, the correlator arises from integral over the full time domain  [^1] The equivalence between in-in formalism and in-out correlation functions has been discussed in [46] for flat space, and [37] for de Sitter correlators. Additionally, here we work in Euclidean signature, the equivalence between Lorentzian de Sitter in-in correlators and Euclidean de Sitter was discussed in [47].
@@
Systematic expansion—Although the factorisation properties of the wavefunction and correlator are closely related, we show that expansions around a given pole are systematically simpler for correlators. In particular, we verify that for the correlator, the first sub-leading term always vanishes. Furthermore, for the total-energy pole, the next two subleading terms take remarkably compact forms that can be generated by acting with a simple differential operator on the ($\ell_{0}$-integrated) amplitude.
@@ II Wavefunction vs correlator
Wavefunction singularities—The wavefunction develops singularities whenever the total energy entering any connected subgraph $\mathfrak{g}\subseteq G$ vanishes [1]. In this context, we use the convention that any edge entering the sub-graph from outside, is replaced by a half-edge attached to the corresponding vertex, with the half-edge inheriting the edge variable of the original. For instance, consider the sub-graph of the two-cycle shown in figure 1 consisting of only the top edge, for which we have:
@@
As an example, the bubble singularities are:
@@
That only poles of the type $\sum_{v}x_{v}+\sum_{e}y_{e}$, and never with $2y_{e}$, is true for any correlator, and is seen directly from the integral representation (4). To see why, all we need is to isolate one particular edge integration, pick one of the terms in $G_{F}$ (e.g.: $\theta(t_{i}-t_{j})$), and isolate the integration regions where the vertex times have opposite sign. One part is:
$\int_{-\infty}^{\infty}dt_{i}dt_{j}e^{-x_{i}t_{i}}e^{-x_{j}t_{j}}e^{-y(t_{i}-t_{j})}\theta(t_{i}-t_{j})\theta(-t_{i})\theta(t_{j})=0\,$, (14)
The other part is:
$\int_{-\infty}^{\infty}dt_{i}dt_{j}e^{-x_{i}t_{i}}e^{-x_{j}t_{j}}e^{-y(t_{i}-t_{j})}\theta(t_{i}-t_{j})\theta(t_{i})\theta(-t_{j})\,$, (15)
@@
This allows us to restrict our analysis of the correlator to graphs in which each pair of vertices is connected by at most one edge.
@@ III Soft Limits
Connected subgraphs—We now explore the case where we take all vertex variables in a graph $G$ to zero, except those in a connected subgraph $\mathfrak{g}$  [^2] Here we are treating site variables, $x_{v}$, to be independent of the edge variables, $y_{e}$, so we can formally take the limit.. Choosing $x_{v_{*}}$, where $v_{*}\in\mathfrak{g}$, as the root in (19), then the integral factorizes into the correlator of $\mathfrak{g}$, and the amplitude of a graph where $\mathfrak{g}$ is contracted to a single vertex, which we denote by $G/\mathfrak{g}$.:
$\langle\widehat{G}\rangle$ $=\langle\hat{\mathfrak{g}}\rangle\times\tilde{\mathcal{A}}({G/\mathfrak{g}}),\;\tilde{\mathcal{A}}({G/\mathfrak{g}})=\int\limits_{-\infty}^{\infty}ds_{e}\prod_{e\in E(G/\mathfrak{g})}\widehat{G}_{F}(s_{e})$
$\langle\hat{\mathfrak{g}}\rangle$ $=\int\limits_{-\infty}^{\infty}dt_{*}ds_{e^{\prime}}e^{-\sum_{v\in\mathfrak{g}}x_{v}|t_{*}+S_{v}|}\prod_{e^{\prime}\in E(\mathfrak{g})}\widehat{G}_{F}(s_{e^{\prime}})]\,.$ (20)
Where $\tilde{\mathcal{A}}(G)$ is simply the scattering amplitude with all zeroth components of external momenta set to zero. When $\mathfrak{g}$ is a single vertex, then the property holds for loop graphs too, and we obtain the $\ell_{0}$-integrated amplitude:
$\langle\widehat{G}\rangle|_{x_{v_{*}}\neq 0}=\frac{2}{x_{v_{*}}}\times\tilde{\mathcal{A}}(G),$ (21)
For a connected graph, the $l_{0}$-integrated amplitude $\mathcal{\tilde{A}}(G)$ satisfies a Berends-Giele type recursion
$\tilde{\mathcal{A}}(G)=\sum_{v\in V(G)}\prod_{\mathfrak{g}\in\kappa(G\setminus v)}\left.\frac{\tilde{\mathcal{A}}(\mathfrak{g})}{\mathcal{E}(\mathfrak{g})}\right|_{x_{v}=0},$ (22)
where the product runs over the connected components of the graph $G\setminus v$ and the $\mathcal{E}(G)$ are the usual tube functions defined in (10) evaluated at $x_{v}=0$. The recursion terminates on graphs consisting of a single vertex, for which we have $\tilde{A}(v)=1$.
@@
$\langle\widehat{G}\rangle=\left(\int\limits_{-\infty}^{\infty}ds_{e^{\prime}}\prod_{e^{\prime}\in E(G/P_{*})}\widehat{G}_{F}(s_{e^{\prime}})\right)\times$
@@
$\langle\widehat{G}\rangle=\left(\prod_{e\in E(G/P_{*})}\frac{2}{y_{e}}\right)\times$
$\sum_{p_{*}\in\{x_{*},x,y_{e}\}}2i\mathop{\mathrm{Res}}_{p\to ip_{*}}\frac{2x_{*}x\prod_{e\in E(P_{*})}2y_{e}}{(p^{2}+x_{*}^{2})(p^{2}+x^{2})\prod_{e\in E(P_{*})}(p^{2}+y_{e}^{2})}.$ (24)
@@ IV Expanding the correlator
We now analyse the simplifications that occur in expanding the correlator about an arbitrary pole.
Tree-level $\mathcal{O}(\mathcal{E}_{G})$-expansion—We begin by examining the expansion around $\epsilon=0$, when we rescale the variables $x_{v}=\epsilon\bar{x}_{v}$ and $\tau=\epsilon t_{*}$, such that we are effectively expanding around the total energy pole. Then, (19) becomes
$\langle\widehat{G}\rangle=\frac{1}{\epsilon}\int_{-\infty}^{\infty}d\tau\int_{-\infty}^{\infty}ds^{E}F(\epsilon)\prod_{e\in E(G)}\widehat{G}_{F}(s_{e}),$ (25)
@@
$\langle\widehat{G}\rangle^{(-1)}=\frac{1}{\epsilon}\int_{-\infty}^{\infty}d\tau F(0)\int_{-\infty}^{\infty}ds^{E}\prod_{e\in E(G)}\widehat{G}_{F}(s_{e}).$ (28)
@@
$\langle\widehat{G}\rangle^{(-1)}=\frac{2}{\sum_{v}\bar{x}_{v}}\prod_{e\in E(G)}\frac{2}{y_{e}}.$ (29)
@@
$\langle\widehat{G}\rangle^{(1)}=\frac{-1}{2\sum_{v}\bar{x}_{v}}\sum_{u\neq v}\bar{x}_{u}\bar{x}_{v}\int_{-\infty}^{\infty}ds^{E}\sum_{e\in P_{u}^{v}}s_{e}^{2}\prod_{vv^{\prime}\in E(G)}\widehat{G}_{F}(s_{e}),$ (31)
@@
$\langle\widehat{G}\rangle^{(1)}=-\sum_{u\neq v}\frac{\bar{x}_{u}\bar{x}_{v}}{\mathcal{E}_{G}}\sum_{e\in P_{u}^{v}}\partial_{y_{e}}^{2}\prod_{e\in E(G)}\frac{2}{y_{e}},$ (32)
where the total energy is given by $\mathcal{E}_{G}=\sum_{v}\bar{x}_{v}$.
@@
For an $n$-loop diagram $n$ variables $s_{e}$ will be constrained in terms of the others. Naturally, the equations (27) still hold true, and we recover the $\ell_{0}$-integrated amplitude in this limit for the $\mathcal{E}_{G}^{-1}$ coefficient, and the $\mathcal{E}_{G}^{0}$ coefficient also vanishes. However, for the $\mathcal{E}_{G}$ there is a slight complication. We find from (47), that the cross terms of the form $s_{e}s_{e^{\prime}}$ no longer vanish, given that the integral is no longer odd. This is simply because the argument of the Feynman propagators will be (33) as explained above. Therefore there is no obvious form for the differential operator acting on the amplitude which gives us the $\mathcal{E}_{G}$ term. Nevertheless, it is still possible to generate this coefficient. We first notice that, because of (33), we have a choice in how we parametrize the sum of $s_{e}$ in (47). This is simply the statement that for loop diagrams there is more than one path between two sites. Then, we can consider a one higher-loop graph where we added an edge between two sites, for this diagram we can parametrize (47) such that the path between these two sites is a single $s_{e}$, and thus we do not have the issue of cross terms not vanishing. Then, for loop graphs the differential operator is simply:
$\langle\widehat{G}\rangle^{(1)}=-\frac{1}{2}\frac{1}{\sum_{v}\bar{x}_{v}}\sum_{u\neq v}\bar{x}_{u}\bar{x}_{v}\partial_{y_{uv}}^{2}\tilde{\mathcal{A}}^{uv}(G)|_{y_{uv}\to 0},$ (34)
@@
$\langle\widehat{G}\rangle^{(2)}=-\frac{1}{\sum_{v}\bar{x}_{v}}\sum_{u\neq v}\bar{x}_{u}\bar{x}_{v}\partial_{y_{uv}}^{3}\tilde{\mathcal{A}}^{uv}(G)|_{y_{uv}\to 0}.$ (36)
Beyond $\mathcal{O}(\mathcal{E}_{G}^{2})$ the coefficients become complicated polynomials of $\bar{x}_{v}$, and it is no longer clear whether we can obtain them by acting with a differential operator on an amplitude.
@@
$x_{v}\rightarrow\epsilon\bar{x}_{v}\text{ for all }v\in V(\mathfrak{g}),$
$y_{e}\rightarrow\epsilon\bar{y}_{e}\text{ for all }e\in H(\mathfrak{g}),$ (37)
@@
$\langle\widehat{G}\rangle^{(-1)}=\frac{\tilde{\mathcal{A}}(\mathfrak{g})}{\mathcal{E}_{\mathfrak{g}}}\prod_{\mathfrak{h}\in\kappa(G\setminus\mathfrak{g})}\langle\mathfrak{h}\rangle,$ (38)
where the product is over all connected components of the graph $G$ with $\mathfrak{g}$ removed. In this limit, $\langle\mathfrak{h}\rangle$ is the correlator of the respective connected component with the edges connected to $\mathfrak{g}$ set to zero.
Furthermore, the first derivative with respect to $\epsilon$ is the same as in (27), therefore the sub-leading term in the series expansion around any pole of the correlator vanishes. This places strong constraints on the structure of the correlator. We have found in tree-level examples up to $4$ vertices—including both the $4$-chain and the star—that if we make an ansatz for the correlator inputing only the poles, demanding only the vanishing of subleading terms fixed $97\%$ of the coefficients.
@@ V $\text{Tr}(\phi^{3})$ correlator
$\langle\widehat{\phi}^{n}\rangle|_{|{\bf k}_{*}|\neq 0}=\frac{2}{|{\bf k}_{*}|}\tilde{\mathcal{A}}_{n},$ (39)
@@
Pole expansion—The interesting structure in the expansions above generalizes for the sum of graphs. To obtain the expansion around a given pole we start by re-scaling every momenta contributing to the pole (either external or internal), $\vec{k}_{i}\to\epsilon\,\vec{k}_{i}$, and then expand around $\epsilon=0$. The leading term will be the usual factorization structure of the correlator, and the sub-leading term will vanish. For the total energy pole the $\mathcal{O}(\mathcal{E}_{\text{tot}}^{1})$ term will also have an interesting structure. We can reformulate our rule to obtain the operator acting on the amplitude to retrieve the $\mathcal{E}_{\text{tot}}$ term. The rule for (32) tells us that for every monomial $x_{i}\,x_{j}$ we should act with the sum over all double derivatives with respect to the every edge on the path from $x_{i}$ to $x_{j}$. To reformulate this as a full correlator rule it is best to think about it in terms of the momentum polygon for an $n$-point process. We know each triangulation of the polygon is dual to a graph, and the internal chords (parametrized by $|k_{ij}|$ of a triangulation are dual to edges (above parametrized by $y_{i}$) of the graph. Therefore, the rule of the $\mathcal{E}_{\text{tot}}$ operator is now that for each monomial $|\vec{k}_{i}|\,|\vec{k}_{j}|$ we take the sum over all the double derivatives with respect to the internal chords crossed when we draw a straight line from the outer edge $|\vec{k}_{i}|$ to the outer edge $|\vec{k}_{j}|$. This allows us to re-write (29) as: